OpenStudy (anonymous):

Suppose you roll a six-sided die two times hoping to get two numbers whose sum is less than 4. What is the sample space? How many favorable outcomes are there?

OpenStudy (anonymous):

12 outcomes?

OpenStudy (anonymous):

no your trying to get two nmbers less than the sum of 4

OpenStudy (anonymous):

1+1=2 2+1=3 So two favorable outcomes

OpenStudy (anonymous):

I mean in total for the two rolls, there's tecnically 12 outcomes. One die has 6 sides, if he used it twice it would be 12?

OpenStudy (anonymous):

Since he wants less than 4 there are 3 possibilities, double that it's 6?

OpenStudy (anonymous):

The favorable outcomes in this situation are when the two dice rolled at the same time are equivalent to less than 4 in their sum. Only 2 pairs of numbers are less than 4 those being 1+1 and 2+1 so you only have 2 favorable outcomes. As far as possible outcomes, I believe it's significantly more than 12 cause if you pair 1 to every number through 6 thats 6 outcomes, and if you do the same to every other number you get 1 less pair each time

OpenStudy (anonymous):

So.. 3 doesn't count? (I'm not good with favorable outcomes) Mind using the formula?

OpenStudy (anonymous):

p(event) - Favorable outcomes/Possible outcomes = ?

OpenStudy (anonymous):

P(Sum<4)-3/36 i think is the total amount of outcomes P=3/36

OpenStudy (anonymous):

Haha whoops logged out, my sister explained it to me and I finally got it. You were right, sorry for being complicated lol.